### v2

# CityBench: Evaluating the Capabilities of Large Language Models for Urban Tasks

Abstract
Recently, large language models (LLMs) with extensive general knowledge and powerful reasoning abilities have seen rapid development and widespread application. A systematic and reliable evaluation of LLMs or vision-language model (VLMs) is a crucial step in applying and developing them for various fields. There have been some early explorations about the usability of LLMs for limited urban tasks, but a systematic and scalable evaluation benchmark is still lacking. The challenge in constructing a systematic evaluation benchmark for urban research lies in the diversity of urban data, the complexity of application scenarios and the highly dynamic nature of the urban environment. In this paper, we design CityBench, an interactive simulator based evaluation platform, as the first systematic benchmark for evaluating the capabilities of LLMs for diverse tasks in urban research. First, we build CityData to integrate the diverse urban data and CitySimu to simulate fine-grained urban dynamics. Based on CityData and CitySimu, we design 8 representative urban tasks in 2 categories of perception-understanding and decision-making as the CityBench. With extensive results from 30 well-known LLMs and VLMs in 13 cities around the world, we find that advanced LLMs and VLMs can achieve competitive performance in diverse urban tasks requiring commonsense and semantic understanding abilities, e.g., understanding the human dynamics and semantic inference of urban images. Meanwhile, they fail to solve the challenging urban tasks requiring professional knowledge and high-level reasoning abilities, e.g., geospatial prediction and traffic control task. Codes are openly accessible via https://github.com/tsinghua-fib-lab/CityBench.
1 Introduction
Recent years, large language models (LLMs) with extensive commonsense and reasoning capabilities have achieved excellent results in various fields (Achiam et al., 2023; Touvron et al., 2023), including programming (Hong et al., 2023), mathematics (Wei et al., 2022), visual intelligence (Liu et al., 2024c) and commonsense reasoning (Suzgun et al., 2022; Mialon et al., 2023). Furthermore, powerful LLMs enable many unimaginable research endeavors to become feasible, e.g., agent (Wang et al., 2024a), embodied intelligence (Reed et al., 2022) and AGI (Feng et al., 2024b). These researchers postulate that LLMs, by acquiring extensive world knowledge and commonsense, hold the key to unlocking promising outcomes in these challenging applications. Many works (Achiam et al., 2023; Gurnee & Tegmark, 2023; Roberts et al., 2023) have demonstrated that LLMs can be regarded as ‘world models’ of our life and they are skilled at solving a wide variety of tasks across multiple fields, while other works (Xiang et al., 2024; Yang et al., 2023; Wang et al., 2024b) indicate that LLMs lack an comprehensive understanding of the real physical world and fail to handle many real-life problems. However, these research efforts have primarily focused on the indoor environment (Puig et al., 2018), while neglecting the outdoor environment, specifically the broader urban environment (Batty et al., 2012; Zheng et al., 2014).
Various works have explored the potential of LLMs in modeling urban space and solving urban tasks. For example, researchers evaluate the potential of LLMs on remote sensing understanding tasks (Kuckreja et al., 2023; Zhang & Wang, 2024) and urban visual tasks (Yan et al., 2024; Hao et al., 2024). Gurnee et al. (Gurnee & Tegmark, 2023) and Roberts et al. (Roberts et al., 2023) evaluate whether LLMs acquire the spatial knowledge of the world, such as cities and coordinates. Manvi et al. (Manvi et al., 2023; 2024) try to extract the geospatial knowledge in LLMs to conduct geospatial indicator prediction tasks (Mai et al., 2023). Besides, researchers also explore how to apply LLMs into the realistic urban applications, e.g., traffic control (Da et al., 2023; Lai et al., 2023), traffic prediction (Liu et al., 2024a), mobility prediction (Wang et al., 2023b), visual language navigation (Schumann et al., 2024) and so on. However, on the one hand, these existing works primarily focus on evaluating the static spatial knowledge of LLMs without considering the environment dynamics and interactivity. On the other hand, most of them only focus on one type of task and one modality of data in the urban space, using small dataset that are not scalable globally. Although there are some existing simulators for urban space such as game simulators (Interactive, 2023) and traffic simulators (Lopez et al., 2018), they cannot be directly applied to support the evaluation and significant amount of adaptation work is required. None of them can support the systematic evaluation of LLMs’ capabilities for diverse tasks in urban research, ranging from understanding and reasoning to decision-making tasks.
Figure 1: The framework of CityBench, which consists of a data collector CityData, an activity simulator CitySimu and 8 diverse urban tasks with different modalities. The evaluation data in the benchmark is collected from 13 cities around the world.
In this paper, we propose CityBench, a comprehensive evaluation platform for assessing the capabilities of LLMs to solve the diverse urban tasks. It covers multiple modalities, supports interactive simulations, and includes data from 13 cities around the world. CityBench consists of three modules: a data module CityData for collecting and processing diverse urban data, a simulation module CitySimu for simulating fine-grained urban dynamics, a evaluation module CityBench for the final evaluation of LLMs and VLMs. In CityData, we first collect three kinds of open urban data: geospaital data from Open Street Map, urban visual data from the Google map and ArcGIS, and human activity data from Foursquare and other websites. Then, we build an efficient simulation engine CitySimu to simulate fine-grained urban dynamics and develop various interfaces for controlling the urban dynamics and sensing the urban environments. Furthermore, based on CitySimu, we design a comprehensive benchmark to evaluate the capability of LLMs and VLMs, covering core research problems from various urban research fields. The benchmark comprises two levels of tasks: perception&understanding tasks and decision-making tasks. In perception&understanding tasks, based on the integrated multi-source data from CitySimu, we introduce street view and satellite image understanding and urban space understanding tasks to evaluate the urban geospatial knowledge of LLMs and VLMs. In decision-making tasks, we apply LLMs and VLMs to interact with CitySimu to complete the urban exploration, visual navigation, mobility prediction and traffic signal control task, which require the comprehensive ability of them. In summary, our contribution are as follows,
  • We develop CityData and CitySimu, an urban data collector and processor designed to support diverse urban tasks and applications, as well as an efficient urban simulator for generating find-grained urban dynamics. They provide ease-to-use APIs for controlling urban dynamics and sensing urban environments.
  • We propose CityBench, a comprehensive evaluation benchmark for evaluating the capability of LLMs and VLMs for urban tasks, which includes 4 geospatial understanding tasks and 4 interactive urban decision-making tasks in 13 cities around the world.
  • Extensive experiments on CityBench with 30 well-known open source and commercial LLMs and VLMs demonstrate the effectiveness of CityBench as evaluation benchmark and also discuss the potential and limitation of applying LLMs and VLMs in urban tasks, ranging from understanding and reasoning to decision-making task. Our results shed light for the further application and development of LLMs and VLMs in the future.
Figure 2: The pipeline of building benchmark, including data collection stage, data integration stage, evaluation generation stage and quality control stage.
2 Related Work
LLM Evaluation for Urban Research Researchers from various urban related fields have conducted extensive evaluations of LLM in urban space from different aspects. Zhang et al (Zhang & Wang, 2024) and Kuckreja et al. (Kuckreja et al., 2023) evaluate the performance of multi-modal LLMs on several remote sensing related tasks. Yang et al. (Yang et al., 2024) propose V-IRL benchmark to evaluate the performance of multi-modal LLMs on street view image related tasks including localization and recognition tasks. Mai et al. (Mai et al., 2023) and Manvi et al. (Manvi et al., 2023) use LLMs to predict social indicators like population and education level. Gurnee et al. (Gurnee & Tegmark, 2023) and Bhandari et al. (Bhandari et al., 2023) try to testify whether LLMs know the coordinates of geospatial entity. Mooney et al. (Mooney et al., 2023) and Deng et al. (Deng et al., 2023) use GIS exams to understand the geospatial skills of LLMs. Roberts et al. (Roberts et al., 2023) design GPT4GEO to evaluate the geospatial capabilities of LLM with limited case studies. Different from these works, we first introduce the interactive simulator based systematic evaluation system for LLMs and VLMs, which covers various data modalities, diverse urban task types and differentiated data from 13 cities around the world.
Interactive Decision-making and Urban Simulator Beyond the above static evaluation, researchers also evaluate the capacity of LLMs in the interactive decision making tasks with customized simulators, e.g., web agent (Liu et al., 2023; Zeng et al., 2023) with web environment and embodied intelligence (Yang et al., 2023) with virtual home (Puig et al., 2018). In the urban domain, Schumann et al. (Schumann et al., 2024) apply LLM to do the visual language navigation task in Touchdown (Chen et al., 2019) and Lai (Lai et al., 2023) apply LLMs as the traffic light controller in CityFlow (Zhang et al., 2019) to manage the road traffic. Besides, Yang et al. (Yang et al., 2024) design V-IRL as the environment of street view image related tasks and propose a global scale virtual intelligence benchmark. These works only evaluate the potential of LLMs in single urban decision-making task and most of their results rely on small-scale datasets in limited regions. Different from them, our work builds on an efficient urban simulator with global scale and supports 4 representative urban decision-making tasks with different modality in one benchmark, including urban exploration, outdoor navigation, mobility prediction and traffic control task.
3 Methods
As presented in Figure 1, CityBench is a simulator based evaluation platform with three core components: CityData for collecting and processing diverse urban data, CitySimu for simulating human dynamics and providing an interactive simulation environment, and CityBench for model evaluation on 8 representative urban tasks with different modalities.
3.1 CityData
To present a complete picture of the city’s geospatial structure, semantic features, and human activities, CityData integrates the following globally available data from multiple sources.
Geospatial Data Geospatial data, represented by maps, is the most fundamental data for describing the urban structure including road networks, points of interest (POIs), areas of interest (AOIs), etc. OpenStreetMap (OSM) ***https://openstreetmap.org/ is most widely used open source map data. However, the raw data provided by OSM cannot support the simulation of urban dynamics directly because the relationship between different elements is incomplete such as the connection between buildings and roads. Therefore, we provide a globally available rule-based map building tool within CityData that reconstructs lanes, lane topology, and building-lane connections based on the raw OSM data. The reconstructed map is used as the geospatial base and simulation input in CitySimu.
Urban Visual Data Street view data and satellite images are two types of globally available urban data that contains rich semantic information, which represents the visual of human. Therefore, CityData also integrates the two types of data, the former obtained via Google Maps API and Baidu Maps API, and the latter using the Esri World Imagery as data source. In the open-source CityData python package, street view data is accessed through spatial location and facing direction, and satellite images are acquired through spatial ranges.
Human Activities Data We use the global Foursquare-checkin (Yang et al., 2016) data and a synthetic global origin-destination data (OD data) (loooffeeeey, 2024) as the proxy of human activities to enable the fine-grained human movement simulation. The Foursquare-checkin dataset (Yang et al., 2016) is a long-term user check-in dataset collected from Foursquare https://foursquare.com/ in approximately 400 cities worldwide. It has been widely used in the community over the past ten years (Chen et al., 2024). Origin-destination data is generated by a diffusion model with population from Worldpophttps://www.worldpop.org/ and satellite image from Esri World Imagery as input. While all the user information are anonymized, we follow the license from Foursquare-checkin (Yang et al., 2016) to protect the public privacy.
Figure 3: The simulation framework of CitySimu, including base environment APIs, interactive objects, simulation APIs and language APIs. Besides, supported task examples also present the relation between simulation APIs and evaluation tasks.
3.2 CitySimu
Building on CityData, CitySimu simulates the urban dynamics and provide diverse easy-to-use APIs for the interactive operation. As shown in Figure 3, CitySimu contains the base environment APIs for obtaining the static information of environment, three simulation APIs for human and vehicle behavior simulations, language APIs to enable the interaction between CitySimu and LLMs.
Individual Mobility Simulation Based on the geospatial data, the individual mobility simulation constructs a simulator that can simulate an agent moving and exploring within the city. Agents can obtain the POIs and roads around them through API provided by CityData, and thus plan and decide the next lane or POI to travel in to update their locations. For the mobility prediction task in the city scale, the available actions are defined as the POIs around the city. For the urban exploration task in the local street scale, the available actions are defined as the nearby lanes.
Urban Visual Environment Simulation To further support the study of urban visual intelligence (Fan et al., 2023), we follow  (Chen et al., 2019; Mirowski et al., 2018) to construct a urban visual environment simulation with real street view images and map data. In the environment, agent can access the panoramic images of its location via APIs and then select the available actions to move along the road to arrive the destination. In the outdoor visual-language instruction navigation task, given the human-like instruction, agent can observe the panoramic images of its location, extract key elements from them and then decide one direction to go. This can be saw as an extension of individual mobility simulation with visual input.
Traffic Simulation In the former two simulations, we only simulate the individual actions without the interaction with others. Here, we introduce microscopic traffic simulation to model the interaction behaviors between vehicles and provide a traffic control environment. The simulator takes the geospatial data reconstructed from OSM within CityData and the travel demand described by the synthetic global OD data as inputs. It simulates the vehicle behaviors through realistic driving simulation models including the intelligent driver model (IDM) (Treiber et al., 2000) as the car-following model and the randomized MOBIL model (Kesting et al., 2007; Feng et al., 2021) as the lane-change model to obtain the dynamics of all vehicles in the city at each second. Through the sensing APIs, LLMs can obtain data about urban dynamics such as junction queue length, vehicle speed, and road average speed. Through the control APIs, LLMs can intervene in the city’s operation, such as modifying traffic signal lights, modifying the speed limit of the road, etc.
3.3 CityBench
Based on CityData and CitySimu, we design a multi-modal urban evaluation benchmark CityBench to evaluate the capability of LLMs and VLMs. In the following section, we first summarize the whole pipeline and then give introduction to each task. Due to the limits of space, detailed templates and prompts for each task can refer to the appendix.
3.3.1 Pipeline
Figure 2 describe the procedure of building evaluation benchmark. As introduced before, CityData works in the data collection and data processing stage and CitySimu works in the data processing stage. We focus on introducing the evaluation generation stage and quality control stage as follows.
In the evaluation generation stage, we use template based methods and LLMs/VLMs based methods to generate the evaluation questions. For example, for the image geolocalization task, the groundth location is already known when collecting, thus we directly design template based question to convert the image geolocalization task into question answer pair. As for the outdoor navigation task, we employ VLM to act as human annotation experts to annotate the data to generate the navigation instruction with additional inputs. In CityBench, the instructions for urban exploration task and the outdoor navigation task are generated by LLM assisted methods. Instructions for other tasks are generated from template based methods.
Due to the hand-craft designs and potential issues of LLMs, we apply a quality control stage to filter and rewrite the generated questions to obtain a high quality evaluation questions. For questions generated from template based methods, we use LLM as data quality expert to filter the low-quality data and use LLM as data rewritter to rewrite the questions with diverse formats and expressions. For questions generated from the LLMs/VLMs based methods, we use LLM/VLM as the agent with additional information to execute the task to verify the quality of generated instructions. If the generated questions are filtered too much, we will return to the evaluation generation stage to generate new questions again. Finally, the authors of this paper also participate in the quality control stage to filter and rewrite the generated data to ensure the quality of whole benchmark.
After the above stages, we produce the evaluation benchmark with 8 urban tasks. Their relations are presented in Figure 4. Details of each task are introduced as follows.
Figure 4: 8 tasks in CityBench with their metrics.
3.3.2 Perception and Understanding Task
Image Geolocalization Image geolocalization task is to predict the precise location of image based on its context. Street view image is regarded as the recording of urban appearance and play an important role in understanding the urban environment and dynamics (Fan et al., 2023). Thus, we query VLMs with street view image and require them to directly generate the location of image. A good VLM should recognize the important objects from the street image and mapping them into the potential locations. Following  (Haas et al., 2023), we define two subtasks for this task: city name inference and precise latitude and longitude coordinates inference.
Geospatial Prediction Geospatial predictions are important for understanding the global sustainable development especially for developing countries, e.g., poverty estimation (Jean et al., 2016) and population density estimation (Tatem, 2017). One of the most widely used solutions is using satellite images with machine learning methods to predict these socioeconomic indicators. In the benchmark, following setting from (Manvi et al., 2023), we query VLMs with a satellite image as context to predict the population density of it. We use population from Worldpop (Tatem, 2017) as the groundtruth.
Infrastructure Inference Besides, we also introduce the infrastructure inference task which means to recognize the urban infrastructures from the satellite images. The groundtruth of this task is extracted from the OSM by matching predefined infrastructure key words within a fixed spatial range. Given the satellite image and a list of all kinds of infrastructures, VLM is required to generate the infrastructure names appeared in the image.
GeoQA for City Elements Beyond understanding the urban space from the visual perspective, we introduce geographic question answer(GeoQA) (Mai et al., 2021) to test whether LLMs comprehends the fundamental spatial elements (Lynch, 1964) in a city from the concept view, such as road, landmarks and boundary. Following (Lynch, 1964; Feng et al., 2024a), we classify the spatial elements into six groups and design problems for each group.
3.3.3 Planning and Decision Making Task
Different from the static evaluation introduced in the last section, we design four interactive decision making tasks to evaluate the capabilities of LLMs in dynamic and partial observed environments which are more challenging and realistic. With the interaction with the CitySimu and dynamic human activities, LLMs need to understand the important mechanisms and regularity in the urban environments to complete the decision-making tasks.
Mobility Prediction As one of the fundamental task for understanding the human behaviors and urban dynamics, mobility prediction task is to predict the next location of user in the next time window with given the past mobility trajectory. Here we use the the global Foursquare checkin data to support the mobility prediction in the simulator. We follow (Wang et al., 2023b) to conduct the mobility prediction task via LLMs.
Outdoor Navigation Outdoor navigation task is widely used in neurocognitive science (Epstein et al., 2017) as the important benchmark for evaluating the spatial cognition of human and models. As one of the most widely-used settings in outdoor navigation task, vision-language navigation task (Schumann et al., 2024; Yang et al., 2024) requires the model to follow the human-annotated language instruction to arrive to the destination with the nearby street view images as additional input. This task requires the VLMs to acquire the ability of urban visual scene understanding, language understanding and decision-making.
Urban Exploration Here, we define a text based street exploration task to evaluate the zero-shot navigation capability of LLMs in a new city without visual input and instructions. Different from the visual language navigation which require model to follow the language instruction and understand the scene via street view image, our urban exploration task require model to explore the region via the local information (e.g. accessed road names) provided by the simulator during action and its intrinsic knowledge of the whole urban space in the city.
Traffic Signal Control Traffic signal control task is one of the widely studied realistic urban decision making task in recent years (Wei et al., 2019). It is challenging for existing methods due to the dynamic traffics and the generalization issues. It is to generate the future traffic signal schedule by considering the current traffic states and the future traffics. Lai et al. (Lai et al., 2023) propose LLMLight to employ LLM as decision-making agent for traffic signal control problem and demonstrate the generalization of LLMs. Following this work, we evaluate the potential of LLMs as agents for multiple-intersections traffic signal control.
4 Benchmark and Experiments
Table 1: Multi-source data for 13 cities around the world. The spatial coverage of these cities can be found in the appendix.
World Cities Visual Data GeoSpatial Data Human Activity Data
Satellite
Image
Street View
Image
Roads PoI/AoIs
OD flow
(>10)
Checkins
Asia Beijing 1764 7482 17043 276090 1905025 21015
Shanghai 5925 4170 33321 57731 845188 33129
Mumbai 638 6025 6296 60245 309147 31521
Tokyo 1120 5514 33174 1146094 969865 1044809
Europe London 1710 4148 14418 83892 1401404 173268
Paris 238 6044 4443 21950 28362 85679
Moscow 1558 5761 9850 28289 979064 836313
Americas NewYork 320 3934 5414 349348 71705 390934
SanFrancisco 345 4473 4171 73777 61367 100249
SaoPaulo 1332 5184 28714 1681735  311830 808754
Africa Nairobi 336 5987 2972  264101 135332 25727
CapeTown 896 5175 5947 151711 525578 11591
Oceania Sydney 1935 5087 21390 141997 438763 54170
4.1 Settings
Model Deployment To facilitate usage of CityBench, we have implemented local deployment support for the majority of LLMs and VLMs using VLMEvalKit (Duan et al., 2024) and vLLM (Kwon et al., 2023). Additionally, we also support evaluation through the APIs of commercial closed-source models, e.g., OpenAI and commercial open-source models, e.g. DeepInfra §§§https://deepinfra.com/.
Baselines We select well-known LLMs and VLMs as baselines. For VLMs, we select LLaVa-NeXT (Liu et al., 2024c), CogVLM-v2 (Wang et al., 2023a), MiniCPM-LLama3-V-2.5 (OpenBMB, 2024), Qwen-VL-plus and GPT4o. For LLMs, we select LLama3-8B, LLama3-70B, Mistral-7B-v0.2 (Jiang et al., 2023), Mixtral-8x22B-v0.1 (Jiang et al., 2024), DeepSeekv2 (Shao et al., 2024), GPT3.5, and GPT4 (Achiam et al., 2023). We also select representative baselines, including GeoCLIP (Vivanco Cepeda et al., 2024) for street view image geolocalization, RSVA (Wang et al., 2022) for infrastructure inference, RemoteCLIP (Yeh et al., 2020; Liu et al., 2024b) for population prediction, LSTPM (Sun et al., 2020) for mobility prediction and MaxPressure (Varaiya, 2013) for traffic signal control task.
Evaluation Metrics We follow the common practice of each task to define the metrics. Metrics and instances for each task are presented in Figure 4 and Table 6 in appendix. For each task with results from 13 cities, we report the mean value of them in Table 2 and Table 3. More detailed results like standard deviation value can be found in the appendix.
Table 2: Performance of 16 widely-used VLMs on four urban visual tasks in CityBench, where ‘City Infer’ and ‘Loc Infer’ denote the city inference task and the geo-coordinates inference task of street view image, ‘Population’ denotes the geospatial prediction task, ‘Infra’ denotes the infrastructure inference task, ‘Navigation’ denotes the outdoor visual-language navigation task.
Tasks Perception & Understanding Decision-making
City Infer Loc Infer Population Infra Navigation
[…]
GeoCLIP 0.340 0.464 - - - - -
RSVA - - - - 0.655 - -
RemoteCLIP - - 1.966 0.368 - - -
Qwen2VL-2B 0.630 0.407 2.478 0.008 0.657 0.020 679.333
InternVL2-2B 0.238 0.380 3.142 -0.841 0.738 0.247 236.088
InternVL2-4B 0.398 0.397 2.501 -0.144 0.735 0.260 272.445
Yi-VL-6B 0.000 0.105 5.471 -3.967 0.816 0.267 429.683
Qwen2VL-7B 0.688 0.522 2.637 -0.112 0.773 0.153 529.549
LLaVA-NeXT-8B 0.267 0.221 3.31 -0.764 0.796 0.207 361.647
MiniCPM-V2.5-8B 0.262 0.223 3.57 -1.054 0.806 0.260 296.427
InternVL2-8B 0.522 0.728 2.806 -0.320 0.806 0.233 223.971
GLM-4v-9B 0.726 0.000 2.769 -0.516 0.857 0.247 444.793
CogVLM2-19B 0.559 0.326 2.75 -0.301 0.726 0.087 596.056
InternVL2-26B 0.429 0.003 2.683 -0.209 0.790 0.180 526.079
Yi-VL-34B 0.251 0.003 2.510 -0.052 0.790 0.253 384.005
LLaVA-NeXT-34B 0.501 0.408 2.61 -0.163 0.804 0.267 274.036
InternVL2-40B 0.574 0.555 2.514 -0.113 0.808 0.213 364.032
Qwen-VL-plus 0.793 0.645 3.14 -1.028 0.454 0.240 377.622
GPT4o 0.862 0.797 2.32 0.122 0.812 0.180 388.582
4.2 Experimental Results
Performance of VLMs on Urban Visual Tasks Main results on urban visual tasks are presented in Table 2. We find three interesting observations: 1) VLMs perform much better than the traditional baselines on the urban visual task, e.g, city inference via street view image and infrastructure inference, which require commonsense and semantic understanding ability of models. 2) But, VLMs perform poorly on tasks requiring precise numerical values, or have a diminished advantage, e.g., the geospatial prediction task on population and precise geo-coordindates prediction task. 3) Most of the VLMs’ performance on all the tasks are not robust, e.g., some VLMs’ performance is close to zero and large model within the same model series not always perform better than the smaller models, e.g., InternVL2-26B performs worse than InternVL2-8B. Finally, we find that the performance variability across different LLMs and VLMs is primarily influenced by the capabilities of the LLMs backbone. For example, in widely used VLMs for urban tasks, Intern2.5-7B consistently outperforms Qwen2-7B and Mistral-7B in most tasks, which can be attributed to Intern2.5-7B’s superior performance in general NLP tasks at the same parameter scale. Similarly, LLaVA-NeXT-8B demonstrates performance comparable to MiniCPM-V2.5-8B, as both models leverage the same LLM backbone, LLaMA3-8B.
Performance of LLMs on Urban Tasks without Visual Input Results on urban task without visual input are presented in Table 3. Compared with the simple urban visual tasks, these urban tasks require more reasoning abilities and professional knowledge of LLMs. From Table 3, we find that LLMs perform much better than the traditional methods in mobility prediction task which is consistent with the results from Wang et al. (Wang et al., 2023b). Besides, different from the results from Lai et al. (Lai et al., 2023), the performance of LLMs on the traffic signal control task seems poor which are far worse than the classic Max-Pressure method. It is noted that our testing scene is the more challenging multi-intersection signal control problem. Finally, as for the GeoQA task for city elements, we find that even the best LLM GPT4-Turbo still has a long way from the performance ceiling. Additionally, we present results from several VLMs for these tasks at the bottom of Table 3. Compared to their original LLM backbones, the performance of VLMs shows some decline. For instance, CogVLM2-19B performs worse than Llama3-8B in both mobility prediction and urban exploration tasks. Therefore, maintaining the general capabilities of LLMs during the training of VLMs should be a key direction to ensure their effectiveness across a wide range of tasks.
Figure 5: Detailed results of LLMs on mobility prediction task(left) for different cities and image geolocalization task(right) for different models.
Error Analysis of LLMs We find that LLMs often display errors such as logic error, format error, invalid action, refusal to answer, and hallucinations. As shown in Figure 6, Llama3-8B exhibited a logical error in its judgment of time in the task mobility prediction. For the urban exploration task, Qwen2-7B refused to choose the option and instead demanded the user to use a navigation service to solve the problem. Intern2.5-7B directly stated that it lacks expertise in this area and needs more information to answer the question. Llama3-8B provided an invalid option in the traffic signal task, rendering CitySimu unable to perform next action. We notice that the most frequent error is Misformatted, and several instances close to 0 in Table 3 are mostly caused by formatting errors. Thus, one of the promising direction is to reduce these error from LLMs to improve their practicality.
Error Analysis of VLMs Urban visual tasks require the model to make decisions directly without going through an explanation process. As a result, as shown in Figure 7, common errors include format errors, invalid actions, refusal to answer and hallucinations. In the image geolocalization task, InternVL2-2B provides a response that do not follow the required format, while Yi-VL-34B gives an irrelevant invalid response. CogVLM2-19B and Yi-VL-34B, in the geospatial prediction and infrastructure inference tasks respectively, repeat the examples provided in the question and refuse to answer the actual question. Due to the response format requirements of the tasks, the most common error made by VLMs in urban visual tasks is misformatted responses.
Geospatial Bias of LLMs and VLMs To further investigate the difference between LLMs and VLMs, we report the detailed results of mobility prediction task and image geolocalization tasks from 13 cities in Figure 5. Based on the above results, we have made several interesting discoveries. First, we find that the performance of different LLMs varies a lot across different cities, no LLM can always perform best in mobility prediction tasks. Second, we find that the performance of VLMs on visual task like image geolocalization task are significantly biased. Most VLMs perform well in major international cities, but poorly in some lesser-known cities (e.g., CapeTown and Nairobi). We provide simple evidence for this phenomenon by analyzing the number of public websites available on Google and Wikipedia. For more details, see Section A.3 in the appendix. The variability in evaluation results demonstrate the necessity of establishing a global evaluation benchmark, and also highlights the potential shortcomings and areas for improvement of LLMs.
Table 3: Performance of LLMs and VLMs on four urban tasks without visual input in CityBench, where ‘Top 1 Acc’ denotes the ‘Top1 Accuracy’ metric, ‘Queue’ denotes the ‘Queue Length’ metric.
Tasks Understanding Planning & Decision-making
GeoQA Mobility Prediction Urban Exploration Traffic Signal
[…]
LSTPM - 0.114 0.086 - - - -
Baselines Fixed-Time - - - - - 57.870 993.333
Max-Pressure - - - - - 36.898 1345.333
Mistral-7B 0.229 0.090 0.087 0.730 5.382 64.120 853.333
Qwen2-7B 0.289 0.142 0.109 0.697 5.889 62.271 880.000
Intern2.5-7B 0.304 0.118 0.102 0.738 5.552 55.121 1047.667
LLama3-8B 0.297 0.130 0.094 0.747 5.304 57.738 1014.333
Gemma2-9B 0.339 0.131 0.120 0.716 5.679 74.475 651.333
LLMs Intern2.5-20B 0.315 0.116 0.098 0.679 6.243 61.229 958.667
Gemma2-27B 0.349 0.145 0.118 0.713 5.733 56.081 1009.333
Qwen2-72B 0.357 0.155 0.135 0.697 5.887 66.924 793.333
LLama3-70B 0.329 0.159 0.130 0.796 4.941 59.338 959.667
Mixtral-8x22B 0.321 0.155 0.136 0.745 5.339 65.682 821.333
DeepSeekV2 0.358 0.126 0.101 0.698 5.739 56.086 1020.333
InternVL2-2B 0.296 0.000 0.000 0.672 6.015 55.725 1012.000
InternVL2-4B 0.304 0.130 0.102 0.674 6.091 74.499 647.667
InternVL2-8B 0.329 0.142 0.102 0.703 5.714 53.196 1069.667
InternVL2-26B 0.310 0.137 0.107 0.694 5.723 57.512 971.667
InternVL2-40B 0.351 0.159 0.121 0.675 6.041 52.459 1087.000
VLMs Qwen2VL-2B 0.293 0.103 0.075 0.643 6.315 56.097 1003.667
Qwen2VL-7B 0.286 0.144 0.102 0.660 6.155 55.885 995.333
MiniCPM-V2.5-8B 0.308 0.124 0.092 0.708 5.643 56.066 1001.000
LLaVA-NeXT-8B 0.313 0.124 0.084 0.688 5.891 56.184 989.333
GLM-4v-9B 0.296 0.133 0.092 0.680 5.979 53.870 1058.000
CogVLM2-19B 0.282 0.026 0.029 0.710 5.905 55.229 1046.667
GPT3.5-Turbo 0.285 0.152 0.113 0.719 5.473 56.219 1022.000
GPT4-Turbo 0.398 0.147 0.125 0.757 5.184 55.761 1022.333
Figure 6: Error case analysis in mobility prediction, urban exploration and traffic signal control tasks.
Figure 7: Error analysis in image geolocalization, geospatial prediction and infrastructure inference tasks.
5 Conclusion
In this paper, we propose CityBench, a systematic evaluation benchmark for LLMs and VLMs in diverse urban tasks. With the data support from CityData and simulation support from CitySimu, we design 8 important urban tasks in 13 cities to constitute the CityBench for evaluating the capabilities of LLMs and VLMs. Extensive experiments present that LLMs and VLMs exhibit exceptional performance in various urban tasks requiring commonsense and semantic understanding, but fail in challenging urban tasks which require professional domain knowledge and precise numeric calculations. The extensive results from CityBench demonstrate the potential the applying LLMs and VLMs in various urban tasks and also shed light for the future research of developing more powerful LLMs and VLMs for urban tasks.
Appendix A Appendix
A.1 Few-Shot Performance of LLMs in CityBench
Here, we present the few-shot performance of several representative LLMs in Beijing in the Table 4. For all text-based tasks, we use 2-shot as the default few-shot method. As shown in the table, the impact of few-shot learning varies across different models and tasks. For instance, few-shot learning improves performance for Gemma-27B in GeoQA but reduces it for Gemma2-9B on the same task. Similarly, it benefits Gemma2 in traffic signal tasks but proves detrimental for Llama3.
Table 4: Few-shot performance of several representative LLMs in Beijing.
Model@Beijing GeoQA Mobility Prediction Urban Exploration Traffic Signal
Accuracy Top1 Acc F1 Succ Rate Steps Throughput Travel Time Queue Length
Gemma2-9B-fewshot 0.306 0.115 0.101 0.728 6.152 1528 2129.21 64.269
Gemma2-9B-zeroshot 0.339 0.131 0.120 0.716 5.679 1448 2214.451 71.729
Gemma2-27B-fewshot 0.359 0.109 0.076 0.696 6.260 2240 1746.89 31.683
Gemma2-27B-zeroshot 0.349 0.145 0.118 0.713 5.733 2187 1762.182 33.322
LLama3-8B-fewshot 0.288 0.095 0.0546 0.692 6.424 1547 2132.233 61.84
LLama3-8B-zeroshot 0.297 0.130 0.094 0.747 5.304 2128 1873.757 40.941
Llama3-70B-fewshot 0.343 0.089 0.0620 0.796 4.941 1810 1962.493 50.541
Llama3-70B-zeroshot 0.329 0.159 0.130 0.74 5.876 2031 1893.517 43.475
A.2 Details of Quality Control
In CityBench, the authors will participate in the quality control process for some tasks, following the automatic quality control stage. Taking the manual checking of GeoQA as an example, the original data from OpenStreetMap contains low-quality information, with missing or incorrect details about AOI, POI, and roads. When this low-quality data is used in the evaluation task, LLMs may become confused and generate meaningless answers. In such cases, the authors review the questions to ensure that the information in the context is meaningful. However, due to time limitations for participants, we can only randomly sample the evaluation cases. For instances from cities and regions where authors are unfamiliar, we filter out low-quality instances. For instances from cities and regions where authors are familiar, we rewrite low-quality instances using external information, such as commercial map services. Finally, if data quality remains unsatisfactory after filtering and rewriting, we will regenerate a certain number of cases to fill the gaps. In fact, while considering the probabilities of filtering, we generate enough candidate instances during the initial generation.
A.3 Additional Results of Geospatial Bias Analysis
Taking three well-performing cities (New York, London, Paris) and three under-performing cities (Shanghai, CapeTown, Nairobi) as examples, Table 5 illustrates the relationship between the performance of street view image localization tasks and the size of the training corpus for LLMs. The training corpus size is approximated by the number of Google search entries and Wikipedia entries for each city. Compared to the well-performing cities, the under-performing cities have significantly smaller ‘training corpora’ in the public websites. Additionally, open-source models exhibit significantly worse performance than commercial models in terms of geospatial bias. This observation provides initial evidence for analyzing geographical bias, but we believe there are more diverse factors contributing to this phenomenon.
Table 5: The performance of street view image localization tasks in different cities and its relationship with the amount of data in the ‘training corpus’ of LLMs, where the amount of data in the training corpus is approximated by the number of Google search entries and Wikipedia entries for each city.
Cities GPT4o MiniCPM-V2.5 Google Search Pages Wikipedia Pages
Shanghai 59% 0.4% 582,000,000 62,495
CapeTown 81% 0 587,000,000 54,724
Nairobi 58.19% 0 205,000,000 14,226
New York 94.39% 80% 7,070,000,000 1,039,264
London 88.80% 43.8% 5,460,000,000 848,381
Paris 92.80% 49% 4,170,000,000 353,500
A.4 Map building tool in CityData
The map building tool enhances open-source map data to support subsequent behavior simulations, encompassing lane topology recovery, relationship recognition, intersection reconstruction, area of interest (AOI) mapping, point of interest (POI) clustering, basic traffic rule generation, and right-of-way construction. The framework of map building tool is presented in Figure 8.
Figure 8: Framework of map building tool.
Table 6: Detailed information of 8 evaluation tasks in CityBench, including data modality, metric and data instances. Task settings across different cities keep consistent.
CityBench Tasks Modality Metrics Instances Images
Perception& Understanding Image Geolocalization Image Acc, Acc@1km/25km 6500 6500
Geospatial Prediction Image $r^{2}$, RMSE 5739 5739
Infrastructure Inference Image Accuracy, Recall 5739 5739
GeoQA for City Elements Text Accuracy 13126 /
Planning& Decision Making Mobility Prediction Text Top1-Acc, F1 6500 /
Urban Exploration Text Steps, Success Rate 650 /
Outdoor Navigation Image Distance, Steps, Success Rate 650 55984
Traffic Signal Control Text Queue Length, Throughput 1hour $\times$ 13 cities /
A.5 Discussion
Limitations. While our platform is based on the public data from various sources, the quality of different data may play a important role in the evaluation results. In the future, we plan to collect more kinds of tasks with global scale groundtruth data to further improve the reliability and representativeness of benchmark.
Ethical considerations and potential societal impact. Our benchmark is designed for enable the global evaluation of LLMs and VLMs for various cities with different cultures and countries. We try our best to improve the ease-of-use and fairness for cities with different development levels. However, due to the limitation of accessed data, the evaluation results for different cites varies a lot. Therefore, the variation in evaluation results caused by data quality may lead to a certain degree of misunderstanding regarding the performance on some urban problems. We call the whole community for attention to this issue to improve the usability of LLMs across different races and countries, promoting fairness and sustainable development of the world.
Develop foundation model for urban domain. Based on the results of our benchmark, we find existing LLMs perform poorly on many urban tasks, even worse than some classic simple baseline algorithms. Developing LLMs tailored for urban domain is urgently necessary. We hope our benchmark can accelerate this development and we look forward to a more comprehensive and robust evaluation framework for urban domain.
A.6 Detailed Results for Each task
Results of visual perception tasks of multi-modal LLMs are presented in Table 7. In general, commercial LLMs like GPT4o performs best in most tasks, achieving the best results in 5 out of 7 metrics. Among the open source models, LLaVA-NeXT-34B performs best with achieving 3 second best results in 7 metrics. Besides, we can find that the performance of open source LLM varies significantly across different cities which indicates significant geospatial bias of open source LLMs. Detailed analysis about this can refer to section 4.2. Results of geospatial knowledge QA task of LLMs are presented in Table 8. As Table 8 shows, commercial LLMs GPT4-Turbo performs best in 4 out of 6 tasks and the left two performs second best. In summary, in terms of geospatial knowledge about urban spaces, the open-source LLMs lag far behind the commercial LLMs.
Table 9 presents the results of mobility prediction task and urban exploration task, the performance difference between LLMs in decision making tasks are much smaller than the last section. It is amazing to observe that open source LLMs (e.g., LLama3-70B) performs better than commercial LLMs in these two tasks. Results of traffic signal control task are presented in Table 11. Here, we report results from three cities with various traffic volume and environment. In general, commercial LLMs like GPT-3.5 or GPT-4 performs better but small open source LLMs can also achieve competitive results in some cases. One interesting direction is to investigate these difference and find out the underlying factors behind them.
Table 7: Results of VLMs on three representative visual understanding tasks. For most metrics, higher values are better, except for RMSE. In the table, bold denotes the best results, underline denotes the second best results.
Image Geolocalization Population Density Infrastructures
Models
City
Accuracy
Accuracy
@1km
Accuracy
@25km
$r^{2}$ RMSE Precision Accuracy
GeoCLIP 0.340$\pm$0.195 0.0254$\pm$0.0498 0.4643$\pm$0.222 - - - -
RemoteCLIP - - - - - 0.126$\pm$0.20 0.655$\pm$0.226
LSMSDelta - - - 0.368$\pm$0.416 1.966$\pm$0.678 - -
Qwen2VL-2B 0.630$\pm$0.411 0.0123$\pm$0.0279 0.407$\pm$0.380 0.008$\pm$0.303 2.478$\pm$0.603 0.101$\pm$0.176 0.657$\pm$0.295
InternVL2-2B 0.238$\pm$0.293 0.008$\pm$0.023 0.380$\pm$0.381 -0.841$\pm$1.421 3.142$\pm$0.492 0.099$\pm$0.194 0.738$\pm$0.236
InternVL2-4B 0.398$\pm$0.383 0.0126$\pm$0.0357 0.397$\pm$0.427 -0.144$\pm$0.840 2.501$\pm$0.532 0.131$\pm$0.241 0.735$\pm$0.253
Yi-VL-6B 0.000 0.0013$\pm$0.0044 0.1050$\pm$0.265 -3.967$\pm$2.142 5.471$\pm$1.22 0.151$\pm$0.216 0.816$\pm$ 0.166
Qwen2VL-7B 0.688$\pm$0.420 0.011$\pm$0.027 0.522$\pm$0.463 -0.112$\pm$0.428 2.637$\pm$0.800 0.145$\pm$0.221 0.773$\pm$0.215
LLaVA-NeXT-8B 0.267$\pm$0.322 0.0097$\pm$0.0269 0.221$\pm$0.337 -0.764$\pm$0.582 3.31$\pm$0.823 0.245$\pm$0.301 0.796$\pm$0.187
MiniCPM-V2.5-8B 0.262$\pm$0.267 0.0061$\pm$0.0124 0.223$\pm$0.270 -1.054$\pm$0.740 3.57$\pm$0.621 0.251$\pm$0.323 0.806$\pm$0.197
InternVL2-8B 0.522$\pm$0.434 0.032$\pm$0.0574 0.728$\pm$0.247 -0.320$\pm$0.596 2.806$\pm$0.68 0.135$\pm$0.217 0.806$\pm$0.187
GLM-4v-9B 0.726$\pm$0.416 0.000 0.000 -0.516$\pm$1.366 2.769$\pm$0.614 0.017$\pm$0.132 0.857$\pm$0.231
CogVLM2-19B 0.559$\pm$0.333 0.0067$\pm$0.0176 0.326$\pm$0.335 -0.301$\pm$0.737 2.75$\pm$0.504 0.287$\pm$0.320 0.726$\pm$0.262
InternVL2-26B 0.429$\pm$0.385 0.001$\pm$0.002 0.003$\pm$0.007 -0.209$\pm$0.546 2.683$\pm$0.626 0.124$\pm$0.216 0.790$\pm$0.178
Yi-VL-34B 0.251$\pm$0.281 0.0006$\pm$0.002 0.003$\pm$0.007 -0.052$\pm$0.487 2.510$\pm$0.697 0.133$\pm$0.219 0.790$\pm$0.182
LLaVA-NeXT-34B 0.501$\pm$0.387 0.0123$\pm$0.0328 0.408$\pm$0.413 -0.163$\pm$0.603 2.61$\pm$0.643 0.262$\pm$0.297 0.804$\pm$0.187
InternVL2-40B 0.574$\pm$0.447 0.013$\pm$0.034 0.555$\pm$0.481 -0.113$\pm$0.689 2.514$\pm$0.604 0.124$\pm$0.206 0.808$\pm$0.161
Qwen-VL-plus 0.793$\pm$0.288 0.0052$\pm$0.0039 0.645$\pm$0.313 -1.028$\pm$1.667 3.14$\pm$1.016 0.203$\pm$0.233 0.454$\pm$0.170
GPT4o 0.862$\pm$0.079 0.0187$\pm$0.0375 0.797$\pm$0.148 0.122$\pm$0.309 2.32$\pm$0.664 0.248$\pm$0.306 0.812$\pm$0.195
Table 8: Results of different LLMs on GeoQA for city, a geospatial knowledge question answering task for urban space. Here, the higher value is better, bold denotes the best results, underline denotes the second best results.
Models Nodes Landmarks Paths Districts Boundary Others
Mistral-7B 0.2286$\pm$0.0167 0.1950$\pm$0.0083 0.2054$\pm$0.0179 0.1511$\pm$0.0150 0.3269$\pm$0.0232 0.2688$\pm$0.0186
Qwen2-7B 0.2214$\pm$0.0174 0.2762$\pm$0.0121 0.2709$\pm$0.0129 0.2906$\pm$0.0371 0.3208$\pm$0.0118 0.3568$\pm$0.0157
Intern2.5-7B 0.2035$\pm$0.0096 0.2946$\pm$0.0167 0.2783$\pm$0.0142 0.2897$\pm$0.0451 0.4296$\pm$0.0447 0.3296$\pm$0.0183
LLama3-8B 0.2322$\pm$0.0150 0.2765$\pm$0.0143 0.3173$\pm$0.0054 0.2627$\pm$0.0163 0.3569$\pm$0.0109 0.3365$\pm$0.0225
Gemma2-9B 0.2605$\pm$0.0180 0.3988$\pm$0.0443 0.2915$\pm$0.0109 0.3277$\pm$0.0571 0.3738$\pm$0.0336 0.3809$\pm$0.0223
Intern2.5-20B 0.2371$\pm$0.0117 0.3792$\pm$0.0409 0.3422$\pm$0.0080 0.2280$\pm$0.0279 0.3808$\pm$0.0329 0.3249$\pm$0.0222
Gemma2-27B 0.2567$\pm$0.0235 0.3569$\pm$0.0308 0.3608$\pm$0.0112 0.3734$\pm$0.0279 0.3781$\pm$0.0229 0.3715$\pm$0.0177
Qwen2-72B 0.3047$\pm$0.0247 0.3758$\pm$0.0376 0.3849$\pm$0.0399 0.3146$\pm$0.0541 0.4135$\pm$0.0135 0.3521$\pm$0.0220
LLama3-70B 0.2646$\pm$0.0076 0.3304$\pm$0.0320 0.4048$\pm$0.0334 0.2532$\pm$0.0429 0.3727$\pm$0.0112 0.3485$\pm$0.0185
Mixtral-8x22B 0.2633$\pm$0.0155 0.3569$\pm$0.0367 0.3652$\pm$0.0201 0.2730$\pm$0.0370 0.3415$\pm$0.0125 0.3239$\pm$0.0241
DeepSeekV2 0.2904$\pm$0.0065 0.3035$\pm$0.0165 0.4133$\pm$0.0208 0.3518$\pm$0.0399 0.4650$\pm$0.0363 0.3264$\pm$0.0232
GPT3.5-Turbo 0.2477$\pm$0.0260 0.2627$\pm$0.0128 0.3222$\pm$0.0255 0.1909$\pm$0.0244 0.3788$\pm$0.0137 0.3050$\pm$0.0244
GPT4-Turbo 0.3429$\pm$0.0095 0.3938$\pm$0.0393 0.5249$\pm$0.0349 0.3054$\pm$0.0412 0.4485$\pm$0.0198 0.3707$\pm$0.0219
Table 9: Results of LLMs on mobility prediction and urban exploration tasks. The arrow next to the metric indicates the direction of improvement, an upward arrow indicates that higher values of the metric are better.
Mobility Prediction Urban Exploration
Models Top1 Accuracy $\uparrow$ F1 $\uparrow$ Success Rate $\uparrow$ Average Steps $\downarrow$
LSTPM 0.114$\pm$0.0013 0.086$\pm$0.001 - -
Mistral-7B 0.090$\pm$0.0255 0.087$\pm$0.031 0.730$\pm$0.134 5.382$\pm$1.982
Qwen2-7B 0.142$\pm$0.0328 0.109$\pm$0.028 0.697$\pm$0.076 5.889$\pm$1.180
Intern2.5-7B 0.118$\pm$0.0361 0.102$\pm$0.033 0.738$\pm$0.064 5.552$\pm$0.947
LLama3-8B 0.130$\pm$0.0320 0.094$\pm$0.030 0.747$\pm$0.140 5.304$\pm$2.059
Gemma2-9B 0.131$\pm$0.0428 0.120$\pm$0.039 0.716$\pm$0.079 5.679$\pm$1.154
Intern2.5-20B 0.116$\pm$0.0379 0.098$\pm$0.031 0.679$\pm$0.077 6.243$\pm$1.149
Gemma2-27B 0.145$\pm$0.0282 0.118$\pm$0.028 0.713$\pm$0.092 5.733$\pm$1.295
Qwen2-72B 0.155$\pm$0.0317 0.135$\pm$0.032 0.697$\pm$0.073 5.887$\pm$1.097
LLama3-70B 0.159$\pm$0.0430 0.130$\pm$0.034 0.796$\pm$0.085 4.941$\pm$1.592
Mixtral-8x22B 0.155$\pm$0.0419 0.136$\pm$0.042 0.745$\pm$0.133 5.339$\pm$1.961
DeepSeekV2 0.126$\pm$0.0520 0.101$\pm$0.044 0.698$\pm$0.119 5.739$\pm$1.812
GPT3.5-Turbo 0.152$\pm$0.0388 0.113$\pm$0.035 0.719$\pm$0.143 5.473$\pm$2.015
GPT4-Turbo 0.147$\pm$0.0357 0.125$\pm$0.035 0.757$\pm$0.091 5.184$\pm$1.520
Table 10: Results of LLMs on traffic signal tasks on three cities with representative traffic volume. We used real-world mobility demands to generate traffic to support the simulation. The traffic signals we actually control in the experiment are only a part of the trajectory path, while the remaining signals are operated according to the default fixed-time scheduler.
Paris Newyork Beijing
Methods
Throughput
Travel
Time
Queue
Length
Throughput
Travel
Time
Queue
Length
Throughput
Travel
Time
Queue
Length
FixedTime 296 2668.313 46.605 532 2689.421 90.982 2152 1799.275 36.023
MaxPressure 655 2113.981 24.342 867 2298.922 70.121 2514 1560.866 16.232
Mistral-7B 290 2656.038 46.795 574 2634.054 87.807 1696 2059.840 57.759
Qwen2-7B 329 2612.115 45.458 539 2640.588 88.757 1772 1979.836 52.597
Intern2.5-7B 339 2603.762 45.411 582 2601.627 85.473 $2$ 1773.167 34.481
LLama3-8B 335 2605.558 45.283 580 2635.365 86.993 2128 1873.757 40.941
Gemma2-9B 116 2766.652 56.236 390 2716.232 95.461 1448 2214.451 71.729
Intern2.5-20B 328 2612.745 45.838 488 2675.345 95.357 2060 1853.243 42.493
Gemma2-27B 334 2609.845 45.209 507 2674.032 89.712 2187 1762.182 33.322
Qwen2-72B 332 2607.622 45.293 528 2664.426 89.279 1520 2115.454 66.201
DeepSeek-67B 331 2614.453 45.528 505 2682.665 89.781 2237 1765.086 33.575
Llama3-70B 332 2607.275 45.424 516 2664.102 89.116 2031 1893.517 43.475
Mixtral-8x22B 334 2614.251 45.473 526 2655.111 88.567 1604 2093.986 63.006
DeepSeekv2 333 2607.169 45.574 495 2682.194 89.914 2233 1754.923 32.771
GPT3.5-Turbo 323 2618.222 45.768 544 2649.214 87.926 2199 1778.963 34.965
GPT4-Turbo 332 2606.985 45.344 495 2670.238 89.706 2240 1746.608 32.234
Table 11: Detailed results of 13 VLMs on outdoor visual-language navigation tasks on Paris, Newyork and Beijing.
Paris Newyork Beijing
Models
Success
Rate
Average
Steps
Average
Distance
Success
Rate
Average
Steps
Average
Distance
Success
Rate
Average
Steps
Average
Distance
Qwen2VL-2B 0.00 0.10 649.527 0.06 0.40 633.234 0.00 0.14 755.240
InternVL2-2B 0.22 1.84 270.809 0.34 2.18 164.118 0.18 2.40 273.337
InternVL2-4B 0.22 1.62 318.246 0.34 2.00 230.406 0.22 2.44 268.684
Yi-VL-6B 0.26 0.94 437.584 0.34 1.60 361.249 0.20 1.66 490.216
Qwen2VL-7B 0.12 0.60 508.745 0.20 1.12 450.643 0.14 1.02 629.261
LLaVA-NeXT-8B 0.18 0.86 468.477 0.32 1.70 293.783 0.12 1.72 322.683
MiniCPM-V2.5-8B 0.22 1.52 349.067 0.34 2.00 293.465 0.22 2.34 246.749
InternVL2-8B 0.16 1.90 243.161 0.34 2.14 222.562 0.20 2.68 206.189
GLM-4v-9B 0.18 0.68 519.489 0.34 1.46 406.329 0.22 1.72 408.560
CogVLM2-19B 0.06 0.30 605.055 0.10 0.56 547.929 0.10 0.88 635.185
InternVL2-26B 0.14 1.04 521.734 0.18 1.12 574.102 0.22 1.78 482.403
Yi-VL-34B 0.22 0.96 463.780 0.34 1.72 288.500 0.20 1.76 399.736
LLaVA-NeXT-34B 0.22 1.60 308.899 0.34 1.90 247.323 0.24 2.52 265.886
InternVL2-40B 0.22 1.52 339.362 0.26 1.40 453.538 0.16 2.30 299.197
Qwen-VL-plus 0.18 0.96 442.616 0.32 1.76 302.899 0.22 2.16 387.353
GPT4o-Mini 0.34 2.02 296.231 0.32 2.22 226.265 0.28 2.46 234.481
GPT4o 0.24 1.28 429.289 0.14 1.30 405.259 0.16 1.92 331.199
A.7 Parameter Settings
In the experiments, the default generation parameters for LLMs are presented in Table 12. Besides, we modify some parameters for some experiments. For street navigation tasks, the temperature is changed to 0.7 for better exploration. For multi-modal LLM, the temperature of CogVLM2 is 0.8 and the temperature of MiniCPM-V-2.5 is 0.7.
Table 12: Generation parameter settings for LLMs during the experiments.
Pameters
Open Source
LLM(DeepInfra)
Commercial
 API
Open Source
MLLM
Commercial API
for MLLM
temperature 0 0 0.2 1.0
top p 1.0 1.0 1.0 1.0(GPT4o), 0.8(Qwen-VL-plus)
max tokens 512 512 512 100
presence penalty 0 0 0 0
frequency penalty 0 0 0 0
repetition penalty 1 1 1 1
Figure 9: Distribution of satellite and street view images in 13 cities.
A.8 Distribution of satellite and street view images in 13 cities
The spatial distribution of sampled satellite and street view images in 13 cities are presented in Figure 9. In each subfigure, the top figure shows the spatial distribution of sampled satellite images, with blue points indicating the actual collection locations and red points representing the sampled locations in the benchmark. Similarly, the bottom figure shows the distribution of sampled street view images, with blue points for collected images and red points for sampled ones.
A.9 Spatial range of 13 Cities
Selected spatial range of 13 cities are presented in Figure 10.
Figure 10: Maps for 13 cities.
A.10 Prompt Examples
Figure 11: Prompt example for geolocalization with street view image.
Figure 12: Prompt example for infrastructure inference with remote sensing image
Figure 13: Prompt example for geospatial prediction with remote sensing image.
Figure 14: Prompt example for traffic signal control.
Figure 15: Prompt example for mobility prediction.
Figure 16: Prompt example for boundary related GeoQA.
Figure 17: Prompt example for districts related GeoQA.
Figure 18: Prompt example for landmark related GeoQA.
Figure 19: Prompt example for node related GeoQA.
Figure 20: Prompt example for other GeoQA.